# Estimating Spatially Resolved Radiation Fields Using Neural Networks

### v1

@@ Abstract
We present an in-depth analysis on how to build and train neural networks to estimate the spatial distribution of scattered radiation fields for radiation protection dosimetry in medical radiation fields, such as those found in Interventional Radiology and Cardiology. Therefore, we present three different synthetically generated datasets with increasing complexity for training, using a Monte-Carlo Simulation application based on Geant4. On those datasets, we evaluate convolutional and fully connected architectures of neural networks to demonstrate which design decisions work well for reconstructing the fluence and spectra distributions over the spatial domain of such radiation fields. All used datasets as well as our training pipeline are published as open source in separate repositories.
@@ 1 Introduction
The spatial distribution of radiation in a volume is of interest for various applications in radiation protection and occupational dosimetry. Knowledge gained from the analysis of irradiation scenarios can be used for optimizing radiation exposure. This is important for work in nuclear power plants, radioactive waste disposal, or in radiological medicine, as Interventional Radiology (IR), in which medical staff performs surgeries under fluoroscopy. The current state of the art relies on the utilization of personal dosemeters, which are sufficient for spatially homogeneous radiation distributions, but inherently misestimate personal doses in inhomogeneous radiation fields.
In the context of IR, medical personnel are exposed to inhomogeneous radiation fields due to their proximity to the patient, which complicates the accurate assessment of individual doses. The reliability of current personal dosimetry methods is contingent upon the uniform distribution of radiation fields. However, this uniformity assumption does not hold in IR settings. To address this issue, previous research has proposed the use of computational dosimetry systems [1], which are designed to monitor the locations and postures of all individuals involved in an IR procedure. In order to assess doses in these exposure scenarios, such systems are making use of currently available methods, that are mainly based on Monte-Carlo Simulations [2].
@@
To advance radiation protection, it is necessary to develop accelerated methods for the simulation of radiation transport. In this research work, we explore the surrogate model approach for MCS acceleration, whereby we learn spatially resolved radiation field distributions from a Monte-Carlo Simulation. Subsequently, the neural networks serve as efficient models for radiation-field reconstruction, providing a real-time capable radiation-transport simulation method. Furthermore, we have generated and made available a variety of datasets for the training of our models and analogous ones. The generation of our datasets was facilitated by the extension and utilization of the RadField3D [5] simulation software which itself is build upon the Geant4 [6] MCS framework to ensure physical correctness, as Geant4 is well established and validated for the targeted use-case.
Our primary innovation is the creation of three datasets of spatially resolved, three-dimensional fluence and spectra distributions for the training and evaluation of neural networks for the reconstruction of radiations fields for IR procedures. We accompany those with concrete neural network implementations and design recommendations based on ablation studies. Additionally, we define several metrics useful for evaluating and comparing spatially resolved radiation field estimators.
@@ 1.1 Monte-Carlo Simulation in Radiation Protection
The technique of using MCS to calculate the radiation transport of complex measurements is already widely spread in various disciplines of dosimetry, from radiation protection to medical physics. Therefore, a set of well-established and tested general-purpose MCS toolkits for radiation transport already exists, such as Geant4 [6], MCNP [7], and EGSnrc [8]. Since these toolkits are meant for general purpose applications ranging from astrophysics over nuclear reactor physics to dosimetry calculations, fine-tuning the MCS parameters for each specific use case is required. There are frameworks built upon these toolkits, such as GATE [9], which is based on Geant4, and it has already been used for assessing doses to medical staff in nuclear medicine by tailoring simulation models on the basis of real tracking of personnel movements [10]. However, as highlighted by Rondon et al., current MCS capabilities are still far from real-time as they require, at best, several tens of seconds if not minutes to be performed. This is particularly challenging in IR, as the calculation of radiation-fluence distributions in the space around the patient during interventions is computationally demanding and can require hours of simulations. However, real-time simulations capabilities are becoming more and more necessary for computational radiation protection dosimetry of staff, as proposed by prior research projects, such as the PODIUM project [1].
Besides the real-time dosimetry, the AR and VR applications for training interventional radiologists and associated personnel demand even faster execution speeds. Nevertheless, especially in AR and VR we can trade absolute accuracy for speed within reasonable bounds. Those applications were already presented various times of the last years, for example in 2012 by an c-arm training extension for VirtX [11] or by the perceptual study of Rainford et al.in 2023 [3]. All of the already implemented systems would benefit from physically correct radiation fields to display during training, but currently those systems are only able to render radiation fields generated by simplified deterministic algorithms due to strict timing requirements. For those applications, achieving frame rates of at least $90-120\,\mathrm{fps}$ requires total rendering times per volume of approximately $8-11\,\mathrm{ms}$ to avoid perceptible visual latency [12].
@@ 1.2 Computational Personal Dosimetry
Real-world measurements in IR scenarios employing Active Personal Dosimeters have already revealed the complexity of contemporary radiation fields and high local dose rates that professionals are exposed to [13]. Those measurements also implied that the definition of a representational location for wearing a dosimeter on the body of the medical staff (whether it is an APD or legal passive dosimeter) is not possible, and therefore, it is likely to both under- or overestimate the received doses by the staff. In the course of the PODIUM project, the researchers developed a camera-based tracking system that was integrated with a MCS application based on the MCNP [7] toolkit. The simulation divided the room into frustums projected from the surface of a sphere placed around the patient at the isocenter of the scene [14, 15]. The evaluation of each event was conducted as it traversed one of the surface segments of the sphere, effectively projecting it along the normal of that surface segment through the scene. The validation of the simulation was based on these frustums, with measurements of APDs worn by medical personnel during real interventions serving as the ground-truth. These interventions involved tracking the locations and postures of staff members.
@@ 1.3 Hybrid Methods and Surrogate Models for Radiation Transport MCS
Due to its notable slowness, as evidenced in the research by Vanhavere & Van Hoey [16], the MCS is ill-suited for real-time or even interactive applications. To address this issue, recent studies have explored the prospect of enhancing its calculation speed. These efforts entail the replacement of the most computationally inefficient components of the radiation transport simulations, which is refereed to as a hybrid approach. Those hybrid methods can be composed of denoising algorithms. The most promising results in this field are delivered by deep learning methods, that use artificial neural networks for transforming the noisy results of an early aborted MCS to those of a fully executed MCS, but with shorter execution times [17, 18, 19]. Although this approach has been the subject of more extensive research, it exhibits a significant disadvantage, as the MCS needs a certain minimum of statistics to be denoise-able which inevitably comes at the cost of a lower processing speed compared to a single neural network. Consequently, it does not meet the requirements for real-time applications especially in the context of VR and AR applications that we mentioned before.
@@
Other approaches focus on the complete replacement of the MCS algorithm, with artificial neural networks. This innovative approach is called surrogate models and its feasibility was already demonstrated in a few studies for different scenarios [20, 21, 22, 23]. Most of these studies were conducted in the context of nuclear power plants, thus those methods aim to predict neutron radiation fields induced by nuclear isotopes. Therefore, those works were heavily focused on modeling shielding and its positioning regarding a static radiation source. For this use case, Wang et al.proposed the use of a residual Fully Connected Neural Network (FCNN) for predicting neutron radiation fields under various shielding configurations [23]. As they could assume a static radiation source, they split the room into regions with similar conditions regarding the radiation field and trained multiple networks for each region. Concurrently, Ye et al.implemented a 3D- Convolutional Neural Network (CNN) for 3D reconstruction in a similar scenario [21]. This 3D-CNN takes the unshielded radiation fields and shielding voxel geometry, convolutes those two channels down to a flat vector which is then rearranged to a 3D radiation field after the shielding process. Another use case for FCNN is demonstrated in the research work of Yisheng et al.[22], where a FCNN is used to predict the spatially resolved photon flux in a room based on the data of a single detector at a known position. The output of the neural network is a low resolution grid with $2000$ voxels. An example of the use of neural networks in the domain of this research work in the context of IR was provided by the research work of M. Villa Arias [20]. Here, a 3D-UNet is used to transform a CT-Scout-Scan to the surrounding radiation field around the patient induced by the CT-scanner.
@@ 2 Materials and Methods
As there is a lack of public available and in this context usable datasets, our first contribution is the generation and provision of three datasets for training spatially resolved doserate predicting neural networks. For those three datasets, we implemented two distinct neural networks architectures, with additional variants for each, in order to analyze the influence of those architectures and specific hyperparameters on the quality of the predictions. For this purpose, we considered two basic architectures: A FCNN architecture inspired by Neural Radiance Fields [24] and a 3D-U-Net [25] architecture which is based on the CNN architecture, but with an extra transposed path for reconstructing all voxels at once.
@@ 2.1 Datasets
We generated and provided three synthetic datasets of radiation fields for different situations with increasing parameter spaces. The datasets were all generated using the RadField3D [5] dataset generator. The resulting datasets, consists of one radiation field per fixed set of parameters. Those fields are stored in the RadFiled3D file format and contain three channels. One for the photons of the direct X-ray beam, one for the resulting scatter fields and one containing the binary voxel occupation of the modeled geometry. Each of the first two channels contains the following information per voxel: The simulated volumetric photon fluence $\Phi^{3}_{\mathrm{\gamma}}$, the volumetric photon energy distribution $p(E_{\mathrm{\gamma}})$ and the statistical error $\epsilon_{rel}$ as described in the research work about the MCS application RadField3D. Additionally, all input parameters, such as the X-ray tubes output spectra or the tubes position and direction, are stored per field.
@@
The first dataset we generated, contains $1250$ radiation fields. Each field has a resolution of 50×50×50 voxels with each voxel having an extent of $2\,\mathrm{cm}$ in every dimension resulting in a total field extent of $1\,\mathrm{m}$ per dimension. The varying parameter in this first dataset is the direction of the beam with $\phi,\theta\in[0,2\pi)$, while the X-ray tubes output is fixed to a radiation quality of H-100 according to ISO 4037-1 [26] and the beam shape is fixed to a cone beam with an opening angle of $\angle_{\mathrm{Beam}}=10\degree$. The distance of the X-ray tube relative to the origin is $2.5\,\mathrm{m}$. At the origin of the field, the headless torso of an male alderson RANDO phantom is placed to act as the scatter object for the radiation beam. This setting is derived from the validation measurements conducted for the RadField3D data generator [5], and, although simplified, it is considered realistic for modeling this first exposure scenario.
@@
The second dataset contains $2156$ radiation fields and extends the previous one by the radiation source spectra as an varying parameter. Here we used random, but realistic, spectra for modeling primary radiation in the context of IR. The range of the parameters was defined on the basis of empirical feedback from medical physicist and radiologists, mainly gained during the PODIUM project. Therefore, we defined an energy range for the tube output energy $E_{\mathrm{tube}}\in[40,125]\,\mathrm{keV}$. Further, we added aluminum with a random thickness $t_{\mathrm{Al}}\in[2.5,7.5]\,\mathrm{mm}$ and, optionally, copper, with a random-thickness $t_{\mathrm{Cu}}\in[0.0,0.9]\,\mathrm{mm}$ for simulating realistic filtering material used to improve the X-ray beam quality. As the anode material, we set tungsten, which is a standard reference material in IR imaging devices, and we used a random anode angle of $\angle_{\mathrm{Anode}}\in[8,12]\degree$. From that parameter space for the radiation source, we sampled sets of parameters and passed them to SpekPy [27] to generate reasonable X-ray spectra for the X-ray beams in this dataset.
@@
In the third and most complex dataset, dynamic X-ray tube distancing is introduced. Therefore, a total of $3779$ distinct radiation fields was generated. For this dataset, we carried over the same base parameter space as described in the preceding datasets, and extended it by sampling the X-ray tube distance from the scenes isocenter as $d_{\mathrm{tube}}\in[35,75]\,\mathrm{cm}$. In order to make the dataset even more realistic in representing typical real-world scenarios, the X-ray beam was collimated to a rectangular shape, with fixed dimensions of $40\,\mathrm{cm}\times 30\,\mathrm{cm}$ in the plane passing through the origin.
In Tab. 1, we have summarized the distribution of each input parameter inside the three datasets and additionally the distribution of the contained fluences by providing the dynamic ranges $\mathrm{DR}_{\mathrm{dB}}$ and the Gini coefficient $G_{\mathrm{Gini}}$, which are defined as:
$\mathrm{DR}_{\mathrm{dB}}=10\log_{10}\Bigl(\frac{\max(\Phi^{3}_{\mathrm{\gamma}})}{\min(\Phi^{3}_{\mathrm{\gamma}})}\Bigr)$
@@
where $\mathrm{DR}_{\mathrm{dB}}$ is used to describe the actual ranges needed for reconstruction and $G_{\mathrm{Gini}}$ expresses the deviation of the present fluence distribution from a uniform distribution. For the calculation of $G_{\mathrm{Gini}}$, the fluences are expected to be sorted in ascending order. A Gini coefficient of $G_{\mathrm{Gini}}=0$ indicates, that the dataset is uniformly distributed, while $G_{\mathrm{Gini}}=1$ indicates, that the dataset contains only one value or category.
@@
Dataset $\overline{\angle}_{\mathrm{Beam}}\pm\sigma$ $\overline{\mathrm{dist}}_{tube}\pm\sigma$ $\overline{\mathrm{kV}}\pm\sigma$ $\overline{\mathrm{kV_{p}}}\pm\sigma$ $\mathrm{DR}_{\mathrm{dB}}$ $G_{\mathrm{Gini}}$
DS-01 $10.0\degree\pm 0.0$ $2.5\,\mathrm{m}\pm 0.0$ $80.7\,\mathrm{kV}\pm 0.0$ $100\,\mathrm{kV}\pm 0.0$ $43.9\,\mathrm{dB}$ $0.790$
DS-02 $10.0\degree\pm 0.0$ $2.5\,\mathrm{m}\pm 0.0$ $48.7\,\mathrm{kV}\pm 9.3$ $82.1\,\mathrm{kV}\pm 24.6$ $53.5\,\mathrm{dB}$ $0.872$
DS-03 - $0.6\,\mathrm{m}\pm 0.1$ $48.7\,\mathrm{kV}\pm 9.3$ $82.1\,\mathrm{kV}\pm 24.6$ $76.0\,\mathrm{dB}$ $0.991$
By observing the differences between our three datasets, one can see, that by reducing the volume of the direct beam inside the radiation field volume, the dynamic range $\mathrm{DR}_{\mathrm{dB}}$ is also drastically increased. In the case of DS-03 when smaller opening angles of the beam and smaller distances of the X-ray tube from the isocenter are used, particle fluences $\Phi^{3}_{\mathrm{\gamma}}$ span over nearly $8$ orders of magnitudes, instead of the spanning of about $5$ orders of magnitudes for the other datasets. Concurrently, the distribution of fluences is more challenging, as the amount of voxels with fluences below $0.1\,\%$ of the maximum fluence rises by a factor of approximately $253$ from $0.3\,\%$ to $76\,\%$.

### v2

@@ Abstract
We present an in-depth analysis on how to build and train neural networks to estimate the spatial distribution of scattered radiation fields for radiation protection dosimetry in medical radiation fields, such as those found in interventional radiology and cardiology. We present three different synthetically generated datasets with increasing complexity for training, using a Monte-Carlo Simulation application based on Geant4. On those datasets, we evaluate convolutional and fully connected architectures of neural networks to demonstrate which design decisions work well for reconstructing the fluence and spectra distributions over the spatial domain of such radiation fields. All our datasets, as well as our training pipeline, are published as open source in separate repositories.
@@ 1 Introduction
The spatial distribution of radiation in a volume is of interest for various applications in radiation protection and occupational dosimetry. Knowledge gained from the analysis of irradiation scenarios can be used for optimizing radiation exposure. This is important for work, for example, in nuclear power plants, radioactive waste disposal, or in radiological medicine, such as Interventional Radiology (IR), in which medical staff performs surgeries under fluoroscopy. The current state of the art relies on the utilization of personal dosemeters, which are sufficient for spatially homogeneous radiation distributions, but inherently misestimate personal doses in inhomogeneous radiation fields.
In the context of IR, medical staff are exposed to inhomogeneous radiation fields due to their proximity to the patient, which complicates the accurate assessment of individual doses. The reliability of current personal dosimetry methods is contingent upon the uniform distribution of radiation fields. However, this assumption of uniformity does not hold in IR settings. To address this issue, previous research has proposed the use of computational dosimetry systems [1], which are designed to monitor the locations and postures of all individuals involved in an IR procedure. In order to assess doses in these exposure scenarios, such systems make use of currently available methods, that are mainly based on Monte-Carlo Simulations [2].
@@
To advance radiation protection, it is necessary to develop accelerated methods for the simulation of radiation transport. In this research work, we explore the surrogate model approach for MCS acceleration, whereby we learn spatially resolved radiation field distributions from a Monte-Carlo Simulation. Subsequently, the neural networks serve as efficient models for radiation-field reconstruction, providing a real-time capable radiation-transport simulation method. Furthermore, we have generated and made available a variety of datasets for the training of our models and analogous ones. The generation of our datasets was facilitated by the extension and utilization of the RadField3D [5] simulation software which itself is built upon the Geant4 [6] MCS framework to ensure physical correctness, as Geant4 is well established and validated for the targeted use-case.
Our primary innovation is the creation of three datasets of spatially resolved, three-dimensional fluence and spectra distributions for the training and evaluation of neural networks for the reconstruction of radiations fields for IR procedures. We accompany these with concrete neural network implementations and design recommendations based on ablation studies. Additionally, we define several metrics useful for evaluating and comparing spatially resolved radiation field estimators.
@@ 1.1 Monte-Carlo Simulation in Radiation Protection
The technique of using MCS to calculate the radiation transport of complex measurements is already widely spread in various disciplines of dosimetry, from radiation protection to medical physics. Therefore, a set of well-established and tested general-purpose MCS toolkits for radiation transport already exists, such as Geant4 [6], MCNP [7], and EGSnrc [8]. As these toolkits are meant for general purpose applications ranging from astrophysics over nuclear reactor physics to dosimetry calculations, fine-tuning the MCS parameters for each specific use case is required. There are frameworks built upon these toolkits, such as GATE [9], which is based on Geant4, and it has already been used for assessing doses to medical staff in nuclear medicine by tailoring simulation models based on the real tracking of personnel movements [10]. However, as highlighted by Rondon et al., current MCS capabilities are still far from real-time as they require, at best, several tens of seconds if not minutes to be performed. This is particularly challenging in IR, as the calculation of radiation-fluence distributions in the space around the patient during interventions is computationally demanding and can require hours of simulations. However, real-time simulation capabilities are becoming increasingly necessary for computational radiation protection dosimetry of staff, as proposed by prior research projects, such as the PODIUM project [1].
Besides the real-time dosimetry, the AR and VR applications for training interventional radiologists and associated personnel demand even faster execution speeds. Nevertheless, especially in AR and VR, we can trade absolute accuracy for speed within reasonable bounds. These applications have already been presented at various times in the last years, for example in 2012 by a c-arm training extension for VirtX [11] or by the perceptual study of Rainford et al.in 2023 [3]. All current implemented systems would benefit from the display of physically correct radiation fields during training, but at the moment these systems are only able to render radiation fields generated by simplified deterministic algorithms due to strict timing requirements. For these applications, achieving frame rates of at least $90$ to $120\,\mathrm{fps}$ requires total rendering times per volume of approximately $8$ to $11\,\mathrm{ms}$ to avoid perceptible visual latency [12].
@@ 1.2 Computational Personal Dosimetry
Real-world measurements in IR scenarios employing Active Personal Dosimeters have already revealed the complexity of contemporary radiation fields and high local dose rates that professionals are exposed to [13]. These measurements also implied that the definition of a representational location for wearing a personal dosemeter on the body of the medical staff is not possible, and therefore, it is likely to both under- or overestimate the received doses by the staff. In the course of the PODIUM project, the researchers developed a camera-based tracking system that was integrated with a MCS application based on the MCNP [7] toolkit. The simulation divided the room into frustums projected from the surface of a sphere placed around the patient at the isocenter of the scene [14, 15]. The evaluation of each event was conducted as it traversed one of the surface segments of the sphere, effectively projecting it along the normal of that surface segment through the scene. The validation of the simulation was based on these frustums, with measurements of APDs worn by medical personnel during real interventions serving as the ground-truth. These interventions involved tracking the locations and postures of staff members.
@@ 1.3 Hybrid Methods and Surrogate Models for Radiation Transport MCS
Due to its notable slowness, as evidenced in the research by Vanhavere & Van Hoey [16], the MCS is ill-suited for real-time or even interactive applications. To address this issue, recent studies have explored the prospect of enhancing its calculation speed. These efforts entail the replacement of the most computationally inefficient components of the radiation transport simulations, which is referred to as a hybrid approach. These hybrid methods can be composed of denoising algorithms. The most promising results in this field are delivered by deep learning methods, that use artificial neural networks for transforming the noisy results of an early aborted MCS to those of a fully executed MCS, but with shorter execution times [17, 18, 19]. Although this approach has been the subject of more extensive research, it exhibits a significant disadvantage, as the MCS needs a certain minimum of statistics to be denoise-able which inevitably comes at the price of a lower processing speed compared to a single neural network. Consequently, it does not meet the requirements for real-time applications especially in the context of VR and AR applications previously mentioned.
@@
Other approaches focus on the complete replacement of the MCS algorithm, with artificial neural networks. This innovative approach is called surrogate models and its feasibility was already demonstrated in a few studies for different scenarios [20, 21, 22, 23]. Most of these studies were conducted in the context of nuclear power plants, thus those methods aim to predict neutron radiation fields induced by nuclear isotopes. Therefore, those works were heavily focused on modeling shielding and its positioning regarding a static radiation source. For this use case, Wang et al.proposed the use of a residual Fully Connected Neural Network (FCNN) for predicting neutron radiation fields under various shielding configurations [23]. As they could assume a static radiation source, they split the room into regions with similar conditions regarding the radiation field and trained multiple networks for each region. Concurrently, Ye et al.implemented a 3D- Convolutional Neural Network (CNN) for 3D reconstruction in a similar scenario [21]. This 3D-CNN takes the unshielded radiation fields and shielding voxel geometry, convolutes those two channels down to a flat vector which is then rearranged as a 3D radiation field after the shielding process. Another use case for FCNN is demonstrated in the research work of Yisheng et al.[22], where a FCNN is used to predict the spatially resolved photon flux in a room based on the data of a single detector at a known position. The output of the neural network is a low resolution grid with $2000$ voxels. An example of the use of neural networks in the domain of this research work in the context of IR was provided by the research work of M. Villa Arias [20]. Here, a 3D-UNet is used to transform a CT-Scout-Scan to the surrounding radiation field around the patient induced by the CT-scanner.
@@ 2 Materials and Methods
As there is a lack of publicly available, and in this context, usable datasets our first contribution is the generation and provision of three datasets for training spatially resolved dose rate predicting neural networks. For these three datasets, we implemented two distinct neural network architectures with additional variants for each, in order to analyze the influence of these architectures and specific hyperparameters on the quality of the predictions. For this purpose, we considered two basic architectures: an FCNN architecture inspired by Neural Radiance Fields [24] and a 3D-U-Net [25] architecture which is based on the CNN architecture, but with an extra transposed path for reconstructing all voxels at once.
@@ 2.1 Datasets
We generated and provided three synthetic datasets of radiation fields for different situations with increasing parameter spaces. The datasets were all generated using the RadField3D [5] dataset generator. The resulting datasets consist of one radiation field per fixed set of parameters. These fields are stored in the RadFiled3D file format and contain three channels. One for the photons of the direct X-ray beam, one for the resulting scattered fields and one containing the binary voxel occupation of the modeled geometry. Each of the first two channels contains the following information per voxel: The simulated volumetric photon fluence $\Phi^{3}_{\mathrm{\gamma}}$, the volumetric photon energy distribution $p(E_{\mathrm{\gamma}})$ and the statistical error $\epsilon_{rel}$ as described in the research work about the MCS application RadField3D. Additionally, all input parameters, such as the X-ray tubes output spectra or the tubes position and direction, are stored per field.
@@
The first dataset we generated, contains $1250$ radiation fields. Each field has a resolution of 50×50×50 voxels with each voxel having an extent of $2\,\mathrm{cm}$ in every dimension resulting in a total field extent of $1\,\mathrm{m}$ per dimension. The varying parameter in this first dataset is the direction of the beam with $\phi,\theta\in[0,2\pi)$, while the X-ray tubes output is fixed to a reference radiation quality of H-100 according to ISO 4037-1 [26] and the beam shape is fixed to a cone beam with an opening angle of $\angle_{\mathrm{Beam}}=10\degree$. The distance from the X-ray tube relative to the origin is $2.5\,\mathrm{m}$. At the origin of the field, the headless torso of a male Alderson RANDO phantom is placed to act as the scatter object for the radiation beam. This setting is derived from the validation measurements conducted for the RadField3D data generator [5], and, although simplified, it is considered realistic for modeling this first exposure scenario.
@@
The second dataset contains $2156$ radiation fields and extends the previous one by the radiation source spectra as a varying parameter. Here we used random, but realistic, spectra for modeling primary radiation in the context of IR. The range of the parameters was defined based on empirical feedback from medical physicist and radiologists, mainly gained during the PODIUM project. Therefore, we defined an energy range for the tube output energy $E_{\mathrm{tube}}\in[40,125]\,\mathrm{keV}$. Further, we added aluminum with a random thickness $t_{\mathrm{Al}}\in[2.5,7.5]\,\mathrm{mm}$ and, optionally, copper, with a random-thickness $t_{\mathrm{Cu}}\in[0.0,0.9]\,\mathrm{mm}$ for simulating realistic filtering material used to improve the X-ray beam quality. As the anode material, we set tungsten, which is a standard reference material in IR imaging devices, and used a random anode angle of $\angle_{\mathrm{Anode}}\in[8,12]\degree$. From that parameter space for the radiation source, we sampled sets of parameters and passed them to SpekPy [27] to generate reasonable X-ray spectra for the X-ray beams in this dataset.
@@
In the third and most complex dataset, dynamic X-ray tube distancing is introduced. A total of $3779$ distinct radiation fields were generated. For this dataset, we carried over the same base parameter space as described in the preceding datasets, and extended it by sampling the X-ray tube distance from the scenes isocenter as $d_{\mathrm{tube}}\in[35,75]\,\mathrm{cm}$. In order to make the dataset even more realistic in representing typical real-world scenarios, the X-ray beam was collimated to a rectangular shape, with fixed dimensions of $40\,\mathrm{cm}\times 30\,\mathrm{cm}$ in the plane passing through the origin.
In Tab. 1, we have summarized the distribution of each input parameter inside the three datasets and additionally the distribution of the contained fluences by providing the dynamic ranges ${DR}_{\mathrm{dB}}$ and the Gini coefficient $G_{\mathrm{Gini}}$, which are defined as:
${DR}_{\mathrm{dB}}=10\log_{10}\Bigl(\frac{\max(\Phi^{3}_{\mathrm{\gamma}})}{\min(\Phi^{3}_{\mathrm{\gamma}})}\Bigr)$
@@
where ${DR}_{\mathrm{dB}}$ is used to describe the actual ranges needed for reconstruction and $G_{\mathrm{Gini}}$ expresses the deviation of the present fluence distribution from a uniform distribution. For the calculation of $G_{\mathrm{Gini}}$, the fluences are expected to be sorted in ascending order. A Gini coefficient of $G_{\mathrm{Gini}}=0$ indicates, that the dataset is uniformly distributed, while $G_{\mathrm{Gini}}=1$ indicates, that the dataset contains only one value or category.
@@
Dataset $\overline{\angle}_{\mathrm{Beam}}\pm\sigma$ $\overline{dist}_{\mathrm{tube}}\pm\sigma$ $\overline{\mathrm{E}}\pm\sigma$ $\overline{\mathrm{keV_{p}}}\pm\sigma$ ${DR}_{\mathrm{dB}}$ $G_{\mathrm{Gini}}$
DS-01 $10.0\degree\pm 0.0$ $2.5\,\mathrm{m}\pm 0.0$ $80.7\,\mathrm{keV}\pm 0.0$ $100.0\,\mathrm{keV}\pm 0.0$ $43.9\,\mathrm{dB}$ $0.790$
DS-02 $10.0\degree\pm 0.0$ $2.5\,\mathrm{m}\pm 0.0$ $48.7\,\mathrm{keV}\pm 9.3$ $82.1\,\mathrm{keV}\pm 24.6$ $53.5\,\mathrm{dB}$ $0.872$
DS-03 - $0.6\,\mathrm{m}\pm 0.1$ $48.7\,\mathrm{keV}\pm 9.3$ $82.1\,\mathrm{keV}\pm 24.6$ $76.0\,\mathrm{dB}$ $0.991$
By observing the differences between our three datasets, one can see, that by reducing the volume of the direct beam inside the radiation field volume, the dynamic range ${DR}_{\mathrm{dB}}$ is drastically increased. In the case of DS-03 when smaller opening angles of the beam and smaller distances of the X-ray tube from the isocenter are used, particle fluences $\Phi^{3}_{\mathrm{\gamma}}$ span over nearly $8$ orders of magnitude, instead of the spanning of about $5$ orders of magnitude for the other datasets. Concurrently, the distribution of fluences is more challenging, as the amount of voxels with fluences below $0.1\,\%$ of the maximum fluence, rises by a factor of approximately $253$ from $0.3\,\%$ to $76\,\%$.